OpenStudy (anonymous):

what is log (base 8) 2= x

OpenStudy (anonymous):

\[\log_82=\log_8{8^{1/3}}={1 \over 3} \log_8 8={1 \over 3}\]

OpenStudy (anonymous):

thanks!!

OpenStudy (amistre64):

change of base: ln(2) ---- = x ln(8)

OpenStudy (anonymous):

It means what number x as a power of 8 is equal to 2 so could be written 8^x = 2 Well you know that 2^3 is 8 8^1/3 is 2 or as we say the third root of 8 is 2

OpenStudy (anonymous):

log_{8} 14^{x-1} =13